Given: unknown nucleus with 70 neutrons. This nucl
Answer #1

Nuclear radisu of an atom of mass number A is r = ro A^1/3
mass number of nickel = 58
radius of nickel = r
radius of unknown atom, r = r'
now given relatioknship of the volumes of two nucleii in question
r'^3 = 2*r^3
let mass number of the unknown nucleus be A'
so, ro^3 * A' = 2* ro ^3 * A = 2 * ro^3 * 58
so A' = 116
number of neutrons = 70
number of protons = 116 - 70 = 46
So Atomic number 46 = Palladium